# Knowledge Sharing and Transfer via Centralized Reward Agent for Multi-Task Reinforcement Learning

### v1

@@ Abstract
Reward shaping is effective in addressing the sparse-reward challenge in reinforcement learning by providing immediate feedback through auxiliary informative rewards. Based on the reward shaping strategy, we propose a novel multi-task reinforcement learning framework, that integrates a centralized reward agent (CRA) and multiple distributed policy agents. The CRA functions as a knowledge pool, which aims to distill knowledge from various tasks and distribute it to individual policy agents to improve learning efficiency. Specifically, the shaped rewards serve as a straightforward metric to encode knowledge. This framework not only enhances knowledge sharing across established tasks but also adapts to new tasks by transferring valuable reward signals. We validate the proposed method on both discrete and continuous domains, demonstrating its robustness in multi-task sparse-reward settings and its effective transferability to unseen tasks.
@@ 1 Introduction
Reinforcement learning (RL) has made significant progress across various domains, like robotics Kober et al., (2013), gaming Lample and Chaplot, (2017), autonomous vehicles Aradi, (2020), signal processing Luo et al., (2024), and large language models Shinn et al., (2023). However, environments with sparse and delayed rewards remain a significant challenge, as the absence of immediate feedback hinders the agent from distinguishing the value of states and leads to aimless exploration (Ladosz et al.,, 2022). Reward Shaping (RS) has been proven to be an effective technique to address this challenge by providing additional dense and informative rewards Sorg et al., 2010b ; Sorg et al., 2010a . Concurrently, multi-task reinforcement learning (MTRL) is becoming increasingly important for its ability to share and transfer knowledge across tasks. In this context, the auxiliary rewards infused with task-specific information in RS offer a straightforward entry to distribute knowledge among different tasks. Therefore, we believe that integrating RS techniques into MTRL is a highly promising and intuitive direction to enhance the efficacy of multi-task system learning.
@@
Presently, numerous MTRL algorithms involving knowledge transfer have been developed. Policy distillation methods identify and combine the commonalities across different policies Rusu et al., (2016); Teh et al., (2017); Parisotto et al., (2016); Xu et al., (2024); representation sharing methods extract and disseminate the common features or gradients among agents Yang et al., (2020); D’Eramo et al., (2020); Sodhani et al., (2021); and parameter sharing methods design architectural modules to reuse parameters or layers across networks Sun et al., (2022); Cheng et al., (2023). Despite their potential, these strategies often encounter challenges such as slow adaptation and response to the transferred knowledge, or suffer from delays in reutilizing and comprehending this information. Therefore, leveraging the RS mechanism, which directly adds a metric to the reward function, offers a promising alternative to address these limitations.
@@
Regarding the RS techniques, not all shaped rewards serve effectively as mediums for knowledge transfer. Specifically, the intrinsic motivation based rewards are typically designed using heuristics to generate task-agnostic signals. Examples include incorporating exploration bonuses Bellemare et al., (2016); Ostrovski et al., (2017); Devidze et al., (2022), rewarding novel states Tang et al., (2017); Burda et al., (2018), and encouraging curiosity-driven behaviors Pathak et al., (2017); Mavor-Parker et al., (2022). Although these approaches broaden extensive exploration, they are not directly related to specific tasks and thus lack transferability. Consequently, we focus on another branch of RS methods, the task-contextual rewards, which automatically learn and encode task-specific information, such as the hidden values, contributions of states, or future-oriented insights, that can be shared effectively across various tasks Ma et al., 2024b ; Mguni et al., (2023); Memarian et al., (2021); Ma et al., 2024a .
@@
To share task-related knowledge in MTRL via RS techniques, and inspired by the ReLara framework Ma et al., 2024b , which integrates an assistant reward agent to densify sparse environmental rewards, we propose the Centralized Reward Agent based MTRL framework (CenRA) [^1] The source code is accessible at: https://github.com/mahaozhe/CenRA. The framework consists of two main components: a centralized reward agent (CRA) and multiple distributed policy agents. Each policy agent individually learns control behaviors within their respective tasks and shares their experiences with the CRA, which are collected from interactions with their environment. The CRA extracts common knowledge from these experiences and learns to generate dense rewards that are encoded with task-specific information. These rewards are then distributed back to the policy agents to augment their original environmental rewards. Additionally, given that different tasks may contribute variably to the MTRL system, we introduce an information synchronization mechanism to further balance the knowledge distribution, considering task similarity and agent learning progress, to ensure a system-wide optimal performance. The main contributions of this paper are summarized as follows:
@@ 2 Related Work
Multi-task reinforcement learning (MTRL) has attracted significant attention recently due to its potential to share knowledge across multiple tasks, thereby improving learning performance Caruana, (1993). We discuss existing MTRL literature from three main directions:
@@
Knowledge Transfer methods focus on identifying and transferring task-relevant features across diverse tasks Zeng et al., (2021). Policy distillation Rusu et al., (2016) is a well-studied approach to extract and share task-specific behaviors or representations that many works are built on: Teh et al., (2017) introduced Distral, which distills a centroid policy from multiple task-policies; Parisotto et al., (2016) developed Actor-Mimic, where a single policy is trained to mimic several expert policies from different tasks; while Yin and Pan, (2017) incorporated hierarchical prioritized experience replay buffer to select and learn multi-task experiences; Hessel et al., (2019) further proposed an adaptation mechanism to equalize the impact of each task in policy distillation. Additionally, Xu et al., (2020) explored the transfer of offline knowledge to train policies, and further leveraged online learning for fine-tuning. Bai et al., (2023) introduced a dual-phase learning approach, optimizing individual policies while correcting them across multiple tasks. Mysore et al., (2022) used separate critics for each task to accompany a single actor to integrate their feedback. These methods mitigate gradient interference to an extent, however, balancing the distribution of knowledge across tasks is crucial. Without a careful trade-off, the performance of the entire system could be compromised.
@@
Representation Sharing methods explore architectural solutions of reusing network modules or representing commonalities to the MTRL problem D’Eramo et al., (2020); Devin et al., (2017); Ma et al., 2024c ; Hong et al., (2021). Sun et al., (2022) used a parameter compositional approach to learn and share a subspace of parameters, allowing policies for various tasks to be interpolated within it. Yang et al., (2020) employed soft modularization to learn foundational policies and utilized a routing network to generate probabilities to combine them. He et al., (2024) introduced the Dynamic Depth Routing framework which dynamically adjusts the use of network modules in response to task difficulty. Sodhani et al., (2021) leveraged task-related metadata to create composable representations. Cheng et al., (2023) and Lan et al., (2023) both incorporated attention mechanisms: the former employed attention-based mixture of experts to capture task relationships, while the latter used Temporal Attention for contrastive learning purposes. Although these methods demonstrate efficacy in learning shared representations, they may struggle to fully capture the complexity of highly diverse tasks. Moreover, adapting shared structures to new tasks typically requires extra design efforts.
@@
Single-Policy Generalization methods learn a single policy to solve multiple tasks simultaneously or continuously, in the absence of information from prior policies or task-specific details, in which case, the primary goal is to enhance the policy’s generalization capabilities. Model-free meta-learning techniques have been proposed to enhance the multi-task generalization Finn et al., (2017). Yang et al., (2017) designed a sharing network structure that allows an agent to learn multiple tasks concurrently. Vuong et al., (2019) introduced a confidence-sharing agent to detect and define shared regions between tasks to support single policy learning. Wan et al., (2020) proposed a transfer learning framework to handle mismatches in state and action spaces. Additionally, several methods focus on overcoming gradient interference to enhance the generalization in various tasks Chen et al., (2018); Yu et al., (2020), while Ammar et al., (2014) developed a consecutive learning policy gradient approach. These methods are efficient in saving computational resources, but the generalization ability of the policy may be constrained when faced with out-of-distribution or previously unseen tasks.
@@ 3 Preliminaries
Markov Decision Process (MDP) fundamentally models sequential decision-making problems under uncertainty. An MDP represents the interaction between an agent and its environment as a tuple $\langle S,A,P,R,\gamma\rangle$, where $S$ is the state space, $A$ is the action space, $P:S\times A\times S\rightarrow[0,1]$ is the probability of transiting from one state to another given an action, $R:S\times A\rightarrow\mathbb{R}$ is the reward function, and $\gamma\in[0,1]$ is the discount factor to modulate the importance of future versus immediate reward.
@@
RL with an Assistant Reward Agent (ReLara) Ma et al., 2024b introduces a dual-agent framework designed to tackle the challenge of sparse rewards in RL. The operational architecture of ReLara is shown in Figure 1. Within this framework, the original agent is termed as policy agent, while an assistant reward agent is integrated to enrich the feedback mechanism by generating dense, informative rewards. The reward agent, trained as a self-contained RL agent, autonomously extracts hidden value information from the environmental states (and potentially the actions of the policy agent) to craft meaningful reward signals. These signals significantly improve learning efficiency by providing immediate and pertinent bonuses.
Figure 1: The ReLara framework Ma et al., 2024b .
@@ 4 Methodology
We extend the ReLara framework Ma et al., 2024b , which utilizes a specialized reward agent to extract environmental information and densify the sparse rewards, to MTRL scenarios. Our proposed framework is termed as Centralized Reward Agent based MTRL (CenRL), which incorporates a centralized reward agent (CRA) to support simultaneous reinforcement learning agents across multiple tasks. A high-level illustration of the CenRA framework is shown in Figure 2. The CRA is responsible for extracting general task-specific knowledge from various tasks and distributing valuable information to the policy agents by reconstructing their reward models. The detailed methodology for knowledge extraction and sharing is introduced in Section 4.1. Furthermore, we recognize there are potential disparities in the information that each task could contribute, which might lead to an imbalance in knowledge distribution. To mitigate this, we introduce an information synchronization mechanism by considering two main factors: the similarity of the tasks and the online learning performance of the policy agents, details given in Section 4.2. Finally, the overall framework of CenRA is presented in Section 4.3.
@@ 4.1.1 Problem Formulation
We consider an MTRL setting comprising $N$ distinct tasks $\{T_{1},T_{2},\dots,T_{N}\}$, all executed within the same type of environment $\mathcal{E}$. We assume that the dimensionality and structure of state space $S$ and action space $A$ remain uniform across tasks, to ensure the centralized reward agent processes consistent inputs. Despite this uniformity, each task may feature unique state sets, goals, and transition dynamics. For instance, a series of mazes with the same size and shape but varying map configurations and objects would satisfy this condition. For the environment $\mathcal{E}$, we denote states by $s\in S$ and actions by $a\in A$. Additionally, for each task $T_{i}$, we denote the transition function as $P_{i}(s^{\prime}|s,a)$ and the reward function as $R_{i}(s,a)$.
The CenRA framework consists of a centralized reward agent $\mathcal{A}^{rwd}$ and multiple policy agents $\{\mathcal{A}^{pol}_{1},\allowbreak\mathcal{A}^{pol}_{2},\dots,\mathcal{A}^{pol}_{N}\}$. Each policy agent $\mathcal{A}^{pol}_{i}$ operates independently to complete its corresponding task $T_{i}$, utilizing appropriate RL algorithms as backbones based on the specific characteristics of the task. For example, implementing DQN Mnih et al., (2015) for discrete control decisions, while TD3 Fujimoto et al., (2018) or SAC Haarnoja et al., 2018a for continuous control tasks. Moreover, the policy of the reward agent $\mathcal{A}^{rwd}$ is denoted as $\pi^{rwd}$, and the internal policy of policy agent $\mathcal{A}^{pol}_{i}$ is denoted as $\pi^{pol}_{i}$.
@@ 4.1.2 Centralized Reward Agent
The CRA $\mathcal{A}^{rwd}$ aims to extract environment-relevant knowledge and distribute it to policy agents by generating additional dense rewards to support their original reward functions. Similar to the ReLara framework Ma et al., 2024b , we model the CRA as a self-contained RL agent, yet, as an extension to ReLara, our CRA is designed to concurrently interact with multiple policy agents and their respective tasks. The CRA’s policy $\pi^{rwd}$ generates continuous rewards given both an environmental state and a policy agent’s behavior. Specifically, $\pi^{rwd}$ maps the Cartesian product of the state space and action space, $S\times A$, to a defined reward space, which constrains the rewards to a range of real numbers, $\mathcal{R}=[R_{min},R_{max}]\subset\mathbb{R}$. For simplicity’s sake, we denote the observation of CRA as $s^{rwd}=(s_{i},a_{i})$, where $s_{i}\sim T_{i}$ and $a_{i}\sim\pi^{pol}_{i}(s_{i})$. To distinguish from the environmental reward, the generated reward is termed as knowledge reward, denoted as $r^{knw}$.
@@
The CRA’s update process involves using these stored transitions to optimize the reward-generating actor $\pi^{rwd}$ and the value estimation critic. The temporal difference error is given by:
$\delta_{t}=r^{env}_{t}+\gamma V^{rwd}(s^{rwd}_{t+1})-V^{rwd}(s^{rwd}_{t})|T_{i},$
where $\tau_{t}=(s^{rwd}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})\sim\mathcal{D}$. The objective function for optimizing the critic module is formulated as:
$J(V^{rwd})=\underset{\tau_{t}\sim\mathcal{D}}{\mathbb{E}}[\delta_{t}^{2}],$
which aims to minimize the squared TD error. Concurrently, the actor module is updated through the following objective function:
$J(\pi^{rwd})\!=\!\underset{\tau_{t}\sim\mathcal{D}}{\mathbb{E}}\Big{[}\underset{r^{knw}_{t}\sim\pi^{rwd}(\cdot|s^{rwd}_{t})}{\mathbb{E}}\big{[}\log\pi^{rwd}(r^{knw}_{t}|s^{rwd}_{t})\cdot\delta_{t}\big{]}\Big{]}.$

### v2

# Centralized Reward Agent for Knowledge Sharing and Transfer in Multi-Task Reinforcement Learning

@@ Abstract
Reward shaping is effective in addressing the sparse-reward challenge in reinforcement learning by providing immediate feedback through auxiliary informative rewards. Based on the reward shaping strategy, we propose a novel multi-task reinforcement learning framework that integrates a centralized reward agent (CRA) and multiple distributed policy agents. The CRA functions as a knowledge pool, which aims to distill knowledge from various tasks and distribute it to individual policy agents to improve learning efficiency. Specifically, the shaped rewards serve as a straightforward metric to encode knowledge. This framework not only enhances knowledge sharing across established tasks but also adapts to new tasks by transferring meaningful reward signals. We validate the proposed method on both discrete and continuous domains, including the representative meta world benchmark, demonstrating its robustness in multi-task sparse-reward settings and its effective transferability to unseen tasks.
@@ 1 Introduction
Reinforcement learning (RL) has made significant progress across various domains, like robotics (Kober et al.,, 2013), gaming (Lample and Chaplot,, 2017), autonomous vehicles (Aradi,, 2020), signal processing (Luo et al.,, 2024), and large language models (Shinn et al.,, 2023; Ouyang et al.,, 2022). However, environments with sparse and delayed rewards remain a significant challenge, as the absence of immediate feedback hinders the agent from distinguishing the value of states and leads to aimless exploration (Ladosz et al.,, 2022). Reward Shaping (RS) has been proven to be an effective technique to address this challenge by providing additional dense and informative rewards (Sorg et al., 2010b, ; Sorg et al., 2010a, ). Concurrently, multi-task reinforcement learning (MTRL) is becoming increasingly important for its ability to share and transfer knowledge across tasks. In this context, the auxiliary rewards infused with task-specific information in RS offer a straightforward entry to distribute knowledge among different tasks. Therefore, we believe that integrating RS techniques into MTRL is a highly promising and intuitive direction to enhance the efficacy of multi-task system learning.
@@
Presently, numerous MTRL algorithms involving knowledge transfer have been developed. Policy distillation methods identify and combine the commonalities across different policies (Rusu et al.,, 2016; Teh et al.,, 2017; Parisotto et al.,, 2016; Xu et al.,, 2024; Ma et al., 2024b, ; Ma et al.,, 2023); representation sharing methods extract and disseminate the common features or gradients among agents (Yang et al.,, 2020; D’Eramo et al.,, 2020; Sodhani et al.,, 2021); and parameter sharing methods design architectural modules to reuse parameters or layers across networks (Sun et al.,, 2022; Cheng et al.,, 2023). Despite their potential, these strategies often encounter challenges such as slow adaptation and response to the transferred knowledge, or suffer from delays in reutilizing and comprehending this information. Therefore, leveraging the RS mechanism, which directly adds a metric to the reward function, offers a compelling alternative to address these limitations.
@@
Regarding the RS techniques, not all shaped rewards serve effectively as the medium for knowledge transfer. Specifically, the intrinsic motivation based rewards are typically designed using heuristics to generate task-agnostic signals. Examples include incorporating exploration bonuses (Bellemare et al.,, 2016; Ostrovski et al.,, 2017; Devidze et al.,, 2022), rewarding novel states (Tang et al.,, 2017; Burda et al.,, 2018), and encouraging curiosity-driven behaviors (Pathak et al.,, 2017; Mavor-Parker et al.,, 2022). Although these approaches broaden extensive exploration, they are not directly related to specific tasks and thus lack transferability. Consequently, we focus on another branch of RS methods, the task-contextual rewards, which automatically learn and encode task-specific information, such as the hidden values, contributions of states, or future-oriented insights, that can be shared effectively across various tasks (Ma et al., 2024a, ; Ma et al., 2025b, ; Ma et al., 2025a, ; Mguni et al.,, 2023; Memarian et al.,, 2021).
@@
To share task-related knowledge in MTRL via RS techniques, and inspired by the ReLara framework (Ma et al., 2024a, ), which integrates an assistant reward agent to densify sparse environmental rewards, we propose the Centralized Reward Agent based MTRL fRAmework (CenRA). The framework consists of two main components: a centralized reward agent (CRA) and multiple distributed policy agents. Each policy agent individually learns control behaviors within their respective tasks and shares their experiences with the CRA. The CRA extracts common knowledge from these experiences and learns to generate dense rewards that are encoded with task-specific information. These rewards are then distributed back to the policy agents to augment their original environmental rewards. Additionally, given that different tasks may contribute variably to the MTRL system, we introduce an information synchronization mechanism to further balance the knowledge distribution, considering task similarity and agent learning progress, to ensure a system-wide optimal performance. The main contributions of this paper are summarized as follows:
@@ 2 Related Work
Multi-task reinforcement learning (MTRL) has attracted significant attention recently due to its potential to share knowledge across multiple tasks, thereby improving learning performance (Caruana,, 1993). We discuss existing MTRL literature from three main directions:
@@
Knowledge Transfer methods focus on identifying and transferring task-relevant features across diverse tasks (Zeng et al.,, 2021). Policy distillation (Rusu et al.,, 2016) is a well-studied approach to extract and share task-specific behaviors or representations that many works are built on: Teh et al., (2017) introduced Distral, which distills a centroid policy from multiple task-policies; Parisotto et al., (2016) developed Actor-Mimic, where a single policy is trained to mimic several expert policies from different tasks; while Yin and Pan, (2017) incorporated hierarchical prioritized experience replay buffer to select and learn multi-task experiences; Hessel et al., (2019) further proposed an adaptation mechanism to equalize the impact of each task in policy distillation. Additionally, Xu et al., (2020) explored the transfer of offline knowledge to train policies, and further leveraged online learning for fine-tuning. Bai et al., (2023) introduced a dual-phase learning approach, optimizing individual policies while correcting them across multiple tasks. Mysore et al., (2022) used separate critics for each task to accompany a single actor to integrate their feedback. These methods mitigate gradient interference to an extent, however, balancing the distribution of knowledge across tasks is crucial. Without a careful trade-off, the performance of the entire system could be compromised.
@@
Representation Sharing methods explore architectural solutions of reusing network modules or representing commonalities to the MTRL problem (D’Eramo et al.,, 2020; Devin et al.,, 2017; Hong et al.,, 2021). Sun et al., (2022) used a parameter compositional approach to learn and share a subspace of parameters, allowing policies for various tasks to be interpolated within it. Yang et al., (2020) employed soft modularization to learn foundational policies and utilized a routing network to generate probabilities to combine them. He et al., (2024) introduced the Dynamic Depth Routing framework, which dynamically adjusts the use of network modules in response to task difficulty. Sodhani et al., (2021) leveraged task-related metadata to create composable representations. Cheng et al., (2023) and Lan et al., (2023) both incorporated attention mechanisms: the former employed attention-based mixture of experts to capture task relationships, while the latter used Temporal Attention for contrastive learning purposes. Although these methods demonstrate efficacy in learning shared representations, they may struggle to fully capture the complexity of highly diverse tasks. Moreover, adapting shared structures to new tasks typically requires extra design efforts.
@@
Single-Policy Generalization methods learn a single policy to solve multiple tasks simultaneously or continuously, in the absence of information from prior policies or task-specific details, in which case, the primary goal is to enhance the policy’s generalization capabilities. Model-free meta-learning techniques have been proposed to enhance the multi-task generalization (Finn et al.,, 2017). Yang et al., (2017) designed a sharing network structure that allows an agent to learn multiple tasks concurrently. Vuong et al., (2019) introduced a confidence-sharing agent to detect and define shared regions between tasks to support single policy learning. Wan et al., (2020) proposed a transfer learning framework to handle mismatches in state and action spaces. Additionally, several methods focus on overcoming gradient interference to enhance the generalization in various tasks (Chen et al.,, 2018; Yu et al., 2020a, ), while Ammar et al., (2014) developed a consecutive learning policy gradient approach. These methods are efficient in saving computational resources, but the generalization ability of the policy may be constrained when faced with out-of-distribution or previously unseen tasks.
@@ 3 Preliminaries
Markov Decision Process (MDP) models sequential decision-making problems under uncertainty. An MDP represents the interaction between an agent and its environment as a tuple $\langle S,A,P,R,\gamma\rangle$, where $S$ is the state space, $A$ is the action space, $P:S\times A\times S\rightarrow[0,1]$ is the probability of transiting from one state to another given an action, $R:S\times A\rightarrow\mathbb{R}$ is the reward function, and $\gamma\in[0,1]$ is the discount factor to modulate the importance of future versus immediate reward.
@@
RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024a, ) introduces a dual-agent framework designed to tackle the challenge of sparse rewards in RL. The operational architecture of ReLara is shown in Figure 1. Within this framework, the original agent is termed as policy agent, while an assistant reward agent is integrated to enrich the feedback mechanism by generating dense, informative rewards. The reward agent, trained as a self-contained RL agent, autonomously extracts hidden value information from the environmental states (and potentially the actions of the policy agent) to craft meaningful reward signals. These signals significantly improve learning efficiency by providing immediate and pertinent bonuses.
Figure 1: The ReLara framework Ma et al., 2024a .
@@ 4 Methodology
We propose Centralized Reward Agent fRAmework (CenRA) for MTRL, which incorporates a centralized reward agent (CRA) to support simultaneous reinforcement learning agents across multiple tasks. A high-level illustration of the CenRA framework is shown in Figure 2. The CRA is responsible for extracting general task-specific knowledge from various tasks and distributing valuable information to the policy agents by reconstructing their reward models. The detailed methodology for knowledge extraction and sharing is introduced in Section 4.1. Furthermore, to mitigate the potential disparities in the information that each task could contribute, which might lead to an imbalance in knowledge distribution, we introduce an information synchronization mechanism by considering two main factors: the similarity of the tasks and the online learning performance of the policy agents, details given in Section 4.2. Finally, the overall framework of CenRA is presented in Section 4.3.
@@ 4.1.1 Problem Formulation
We consider an MTRL setting comprising $N$ distinct tasks $\{T_{1},T_{2},\dots,T_{N}\}$, all executed within the same type of environment $\mathcal{E}$. We assume that the shape of state $s\in S$ and action $a\in A$ remain uniform across tasks, to ensure the CRA processes consistent inputs. Despite this uniformity, each task may feature different state spaces, action spaces, goals, and transition dynamics. For instance, a series of mazes with the same size but varying map configurations would satisfy this condition. For each task $T_{i}$, we denote the transition function as $P_{i}(s^{\prime}|s,a)$ and the reward function as $R_{i}(s,a)$.
The centralized reward agent (CRA) is denoted as $\mathcal{A}^{rwd}$ and multiple policy agents are denoted as $\{\mathcal{A}^{pol}_{1},\allowbreak\mathcal{A}^{pol}_{2},\dots,\mathcal{A}^{pol}_{N}\}$. Each policy agent $\mathcal{A}^{pol}_{i}$ operates independently to complete its corresponding task $T_{i}$, utilizing appropriate RL algorithms as backbones. For example, implementing DQN (Mnih et al.,, 2015) for discrete control tasks, while TD3 (Fujimoto et al.,, 2018) or SAC (Haarnoja et al., 2018a, ) for continuous control tasks. Moreover, the policy of CRA $\mathcal{A}^{rwd}$ is $\pi^{rwd}$, and the internal policy of policy agent $\mathcal{A}^{pol}_{i}$ is $\pi^{pol}_{i}$, respectively.
@@ 4.1.2 Centralized Reward Agent
The CRA $\mathcal{A}^{rwd}$ aims to extract environment-relevant knowledge and distribute it to policy agents by generating additional dense rewards to support their original reward functions. Similar to the ReLara framework Ma et al., 2024a , we model the CRA as a self-contained RL agent, yet, as an extension to ReLara, our CRA is designed to concurrently interact with multiple policy agents and their respective tasks. The CRA’s policy $\pi^{rwd}$ generates continuous rewards given both an environmental state and a policy agent’s behavior. Specifically, $\pi^{rwd}$ maps the Cartesian product of the state space and action space, $S\times A$, to a defined reward space, which constrains the rewards to a range of real numbers, $\mathcal{R}=[R_{min},R_{max}]\subset\mathbb{R}$. For simplicity’s sake, we denote the observation of CRA as $s^{rwd}=(s_{i},a_{i})$, where $s_{i}\sim T_{i}$ and $a_{i}\sim\pi^{pol}_{i}(s_{i})$. To distinguish from the environmental reward, the generated reward is termed as knowledge reward, denoted as $r^{knw}$.
@@
The CRA’s update process involves using these stored transitions to optimize the reward-generating actor $\pi^{rwd}$ and the value estimation critic. The objective function for the critic module is:
$J(V^{rwd})=\underset{\tau_{t}\sim\mathcal{D}}{\mathbb{E}}[\delta_{t}^{2}],\quad\delta_{t}=r^{env}_{t}+\gamma V^{rwd}(s^{rwd}_{t+1})-V^{rwd}(s^{rwd}_{t})|T_{i},$ (1)
where $\tau_{t}=(s^{rwd}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})\sim\mathcal{D}$. Concurrently, the actor module is updated through the following objective function:
$J(\pi^{rwd})=\underset{\tau_{t}\sim\mathcal{D}}{\mathbb{E}}\Big{[}\underset{r^{knw}_{t}\sim\pi^{rwd}(\cdot|s^{rwd}_{t})}{\mathbb{E}}\big{[}\log\pi^{rwd}(r^{knw}_{t}|s^{rwd}_{t})\cdot\delta_{t}\big{]}\Big{]}.$ (2)